The length of rectangle is 5 cm longer than its width and its area is 66 cm2. Find perimeter of the rectangle
Answer by mananth(16946)

width be x
length = (x+5)
Area = L * W
x(x+5)=66
x^2+5x=66
x^2+5x-66=0
x^2+11x-6x-66=0
x(x+11)-6(x+11)=0
(x+11)(x-6)=0
x=-11 OR x=6
take the positive value
width = 6 cm
and length = x+5 ==>6+5=11cm